OpenStudy (anonymous):

solve the logarithmic equation log(x+3)+logx=1

OpenStudy (anonymous):

\[\log \left( x+3 \right)+logx=1,\] \[logx \left( x+3 \right)=1=\log10\] \[x \left( x+3 \right)=10\] \[x ^{2}+3x-10=0\] make factors and solve for x.

OpenStudy (anonymous):

-5,2

OpenStudy (anonymous):

x=-5 rejected because domain of log is (0,infinity)